OpenStudy (anonymous):

Suppose 5 men and 7 women are on a crowded elevator. At the next floor, four people get off the elevator. Find the probability that none are women.

OpenStudy (anonymous):

i think it's Pr= n(E)/n(U) = C(5,4)/(2^4) = 0.3125

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

no problem...medal pls :)